OpenStudy (anonymous):

Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 8 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). A)x = 3 B)x = 4 C)x = 5 D)x = 6

OpenStudy (anonymous):

@butterflydreamer @chrisdbest @brebre5564

OpenStudy (anonymous):

@CaptainLlama49

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Lemme see

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

its been a while since ive done this lol do you know what f(x) stands for cx

OpenStudy (anonymous):

NOP

OpenStudy (anonymous):

ARE YOU THERE?

OpenStudy (anonymous):

@CaptainLlama49

OpenStudy (anonymous):

Sorry lol cant help you there

OpenStudy (anonymous):

@Study_together

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (jtvatsim):

Hmm... If I'm reading this right, it looks like they just want you to plug in different x values (from 1 to 6) into the equations.

OpenStudy (jtvatsim):

So, if we started with the "f(x)" equation we are thinking about \[-x^2 + 4x + 12\] and we want to plug in different x values to see what we get. What would you get if you plug in x =1?

OpenStudy (anonymous):

the answer its 6?

OpenStudy (jtvatsim):

Well, let's see. What did you do to get that answer?

OpenStudy (anonymous):

i calculate: −x2+4x+12

OpenStudy (anonymous):

\[f(x) = −x^2 + 4x + 12 \]\[g(x) = x + 8 \]\[f(x)=g(x)\iff -x^2+4x+12=x+8\]

OpenStudy (jtvatsim):

OK, that sounds right. Maybe there was just a calculation error. I did -(1)^2 + 4(1) + 12 = -1 + 4 + 12 = 15. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, so in the "f(x)" row under the "1" column you would fill in 15.

OpenStudy (jtvatsim):

You should have something like this so far 1 2 3 ..... f(x) 15 g(x)

OpenStudy (jtvatsim):

It's a little tedious, but we have to repeat our calculations for all x's up to 6. Next is x = 2. Plug that into -x^2 + 4x + 12 and see what you get. I'll check your answer. :)

OpenStudy (anonymous):

i get 6? i dont know if i'm right

OpenStudy (jtvatsim):

OK, can you write your steps like I did before?

OpenStudy (anonymous):

ok let me see

OpenStudy (jtvatsim):

sure, take your time. :)

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

f(x)=−x2+4x+12 g(x)=x+8 f(x)=g(x)⟺−x2+4x+12=x+8

OpenStudy (jtvatsim):

Well, not exactly. But good job copying all those formulas! :) I meant when you plug in x = 2 into -x2+4x+12. Like this -(2)^2 + 4(2) + 12 = ... What comes next?

OpenStudy (jtvatsim):

so, -(2)^2 = -4 and 4(2) = 8 so we get -(2)^2 + 4(2) + 12 = -4 + 8 + 12 = ... could you finish this part?

OpenStudy (anonymous):

16?

OpenStudy (jtvatsim):

Yes! Exactly! :D

OpenStudy (jtvatsim):

Good job! That's all I meant by "show your steps". Anyways, we now have the second piece of our table 1 2 3 f(x) 15 16 g(x) Does this make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, let's keep going. Next is x = 3. Let's plug this into -x^2 + 4x + 12 and see what we get. I'll help get you started: -(3)^2 + 4(3) + 12 = What comes next?

OpenStudy (anonymous):

=15?

OpenStudy (jtvatsim):

Good!

OpenStudy (anonymous):

so, whats nexts?

OpenStudy (jtvatsim):

Alright, so we still need to do the same thing for x = 4, x = 5, and x = 6. Then we will be done plugging into the -x^2 + 4x + 12 equation. See if you can do these: -(4)^2 + 4(4) + 12 = -(5)^2 + 4(5) + 12 = -(6)^2 + 4(6) + 12 =

OpenStudy (anonymous):

yea give a second

OpenStudy (jtvatsim):

Of course, that's a lot of numbers. :)

OpenStudy (anonymous):

.12 .7 .0

OpenStudy (jtvatsim):

Very good! You got them all!

OpenStudy (jtvatsim):

OK, phew... almost done. We have finally completed the first row of the table: 1 2 3 4 5 6 f(x) 15 16 15 12 7 0 g(x)

OpenStudy (anonymous):

so whats next's?

OpenStudy (jtvatsim):

Next is the easier formula. We do the same thing we did for f(x), but now for g(x). Plug in x = 1, x = 2, x = 3, x = 4, x = 5, and x = 6 into x + 8 (that's g(x)).

OpenStudy (jtvatsim):

As long as we don't get freaked out, we can see that this is pretty quick to figure out. 1 + 8 = 2 + 8 = 3 + 8 = 4 + 8 = 5 + 8 = 6 + 8 =

OpenStudy (jtvatsim):

These answers will go into the second row of our table.

OpenStudy (anonymous):

ok let me solve

OpenStudy (anonymous):

.9 .10 .11 .12 .13 .14

OpenStudy (jtvatsim):

Great job! You've got it!

OpenStudy (jtvatsim):

Let's put these into the table: 1 2 3 4 5 6 f(x) 15 16 15 12 7 0 g(x) 9 10 11 12 13 14

OpenStudy (jtvatsim):

Cool! The table is finished like the problem wanted. Their last question was, when does f(x) = g(x) ? We just have to look at the table and see when the two numbers are the same. Which x value seems to have the same number for both f(x) and g(x)?

OpenStudy (anonymous):

yea, f(x) and g(x) have the 12

OpenStudy (jtvatsim):

Good! There's the match, so x = ? to make this happen. Our choices are: A)x = 3 B)x = 4 C)x = 5 D)x = 6 What do you think?

OpenStudy (anonymous):

4?

OpenStudy (jtvatsim):

Yep! That's the one! If we weren't sure we could just confirm that -(4)^2 + 4(4) + 12 = 12 and 4 + 8 = 12 just like we wanted. Finally! We are done! Congrats!

OpenStudy (anonymous):

thank you so much! you explain good! can you help me with 3 more please?

OpenStudy (jtvatsim):

I wish I could, but I've got to go for now, I've got some of my own homework to finish. Sorry. :( Good luck! I'm sure someone else can help you out. :)

OpenStudy (anonymous):

its fine

OpenStudy (jtvatsim):

Have a good night, or morning, or whatever time it is where you are. :)

OpenStudy (anonymous):

night